OpenStudy (anonymous):

The coordinates of the point have equal absolute value, the point is in the second quadrant, and the distance of the point from the origin is 2. x=????? y=????

OpenStudy (anonymous):

Well, it means that the y and x have the same number but opposite signs. Also, you know that the distance is 2 so that is the hypotenuse of a triangle with x and y as the legs. It is in the second quadrant so x is negative and y is positive. This means that the actual values are x=-sqrt(2) and y=sqrt(2).

OpenStudy (anonymous):

Because sqrt(2)^2 = 2 and 2+2 = 2^2 which satisfies the pythagorean theorem

OpenStudy (anonymous):

ok awesome, that worked! Thanks that makes more sense

OpenStudy (anonymous):

Glad to help